Circular No. 5967

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET RUSSELL 2 (1994e) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery of

this comet with the Spacewatch reflector at Kitt Peak. The following positions are in close agreement with the prediction on MPC 18259 (ephemeris on MPC 23142):

1994 UT  α_{2000}  δ_{2000}  m_{1}
Apr. 5.47858  16^{h}16^{m}33.^{s}67  −23^{o}25′05′′.1  21.1
5.48701  16 16 33.69  −23 25 07.4  21.3
5.49579  16 16 33.70  −23 25 11.3  21.3
6.48003  16 16 37.55  −23 30 16.2  21.2
6.48800  16 16 37.61  −23 30 18.4  21.5
6.49628  16 16 37.54  −23 30 21.3  22.0

GRO J0422+32 E. Martin and M. Serra-Ricart, Instituto de Astrofisica de Canarias

(IAC), La Laguna, communicate: “After successive outbursts of decreasing intensity (cf. IAUC 5921) the x-ray transient GRO J0422+32 has been observed to be in a quiescent state since Feb. 1 via the IAC 0.80-m telescope of Teide Observatory. Photometric CCD monitoring on Mar. 14 and 15 reveals modulation of amplitude 0.3 ± 0.1 mag in the I band and a shape suggesting ellipsoidal variations of period 2.5 hr. These data are consistent with a secondary’s orbital period of 5.1 hr, the shortest known among x-ray transients. Further CCD observations taken with the 2.5-m Nordic Optical Telescope at La Palma give the following photometry: Mar. 26.90 UT, I = 18.87 ± 0.04, (R − I)_{c} = +0.96 ± 0.08, (V − I)_{c} = +1.8 ± 0.1. These colors suggest that the secondary is a K-type star.”

SUPERNOVA 1994I IN NGC 5194 Further magnitude estimates: Mar. 31.58 UT, 14 (S. Sasaki, Hazaki,

Japan; T-Max 400 film; independent discovery); 31.95, 14–15 (M. Moeller and C. Pawellek, Timmendorfer Strand, Germany; prediscovery trailed image with unfiltered CCD); Apr. 1.997, 13.3 (J. Bustamante-Donas, Madrid, Spain; prediscovery unfiltered CCD image); 2.57, 14 (A. Yonezawa, Outoumura, Japan; Fuji G800 film; independent discovery); 7.16, 13.3 (W. G. Dillon, Missouri City, TX; visual).

Corrigendum. On IAUC 5966, line 12, for P. Berland read P. Berlind

1994 April 7 Daniel W. E. Green